OpenStudy (anonymous):

find the are of the triangle. also find the length of the missing side. 40 on the right, 35 on the bottom. please show all work.

OpenStudy (anonymous):

is it a right anged triangle....

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

area=1/2 *b*h area =1/2 *40 *35 area =700 unit square

OpenStudy (anonymous):

got it........

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

welcum